4. A simple linear regression was fit for a dataset with 35 data points. The sample variance of the response was found to be 2.4 and 62 was found to be 1.3. What was the value of R2 for this data?
Answer #1

here as SST =(n-1)*sample variance =34*2.4=81.6

and SSE =(n-2)*pooled variance =33*1.3=42.9

hence R2 =1-SSE/SST =1-42.9/81.6=0.4743